# Not all Chess960 positions are equally complex

### v2

@@ II.2.1 Asymmetry in Decision Complexity
which quantifies the differential complexity: positive values indicate Black faces more difficult decisions, while negative values indicate White’s burden is greater. Analysis of all 960 starting positions (Fig. 4(d)) reveals a mean asymmetry of $\langle A\rangle=-0.25$ bits (and standard deviation $\pm 0.69$ bits), suggesting a weak but systematic bias wherein White generally faces slightly more complex decision trees. This negative mean may reflect the fundamental asymmetry inherent in chess: White moves first and must immediately confront the full complexity of opening choices, while Black’s initial responses are partially constrained by White’s play. The substantial intra-position standard deviation of $1.42$ bits (from 2.51-2.51 to +1.84+1.84 bits) indicates high game-to-game variability, reflecting the sensitivity of decision complexity to specific move sequences.
II.2.2 Complexity and time to make a move
If the quantities $\Delta$ and $S(10)$ are meaningful indicators of the intrinsic complexity of a position, one naturally expects them to correlate with the time players spend deciding on a move. Intuitively, positions with a small $\Delta$ (i.e., many nearly equivalent candidate moves) or a large entropy $S(10)$ (i.e., a broad distribution of plausible continuations) should require more cognitive effort, and therefore more time, to resolve [ref referee].
To test this hypothesis, we analyze a freely available database of online games played on Lichess with move-level time stamps [28]. For each position, we compute $\Delta$ and $S(10)$ using Stockfish at depth 15, and measure the mean time $\tau$ required by players to make a move. We then study the dependence of $\tau$ on $\Delta$ and $S(10)$ (Fig. 5). The dataset consists of $1000$ games, corresponding to a total of 70 15070\,150 moves.
Figure 5: (a) Mean thinking time $\tau$ as a function of $\Delta$. (b) Mean thinking time $\tau$ as a function of the entropy $S(10)$. Data are computed from $1000$ games (total of 70 15070\,150 moves) from [28], using Stockfish at depth 15.
Since players of different strengths perceive and handle complexity differently, we account for Elo-dependent effects by adapting the reference threshold $\Delta_{0}$ as a function of the Elo rating. Specifically, we use
$\Delta_{0}=\begin{cases}100,&\text{Elo}<1400,\\[4.0pt] 100-90\,\frac{\text{Elo}-1400}{1000},&1400\leq\text{Elo}\leq 2400,\\[4.0pt] 10,&\text{Elo}>2400.\end{cases}$ (6)
As expected, we observe a clear decrease of the average thinking time $\tau$ with increasing $\Delta$, and a corresponding increase of $\tau$ with $S(10)$. Deviations from these trends are visible for very small $\Delta$ or very large $S(10)$. These fluctuations can be attributed to additional factors not captured by our complexity measures, such as opening preparation: opening positions often exhibit high complexity and small $\Delta$, yet moves are played rapidly due to memorization.
Despite these effects, the overall trends provide strong empirical support for $\Delta$ and $S(10)$ as relevant proxies for decision-making complexity in chess. They capture, in a quantitative manner, the intuitive notion that positions with many competing continuations or poorly differentiated evaluations are cognitively more demanding and require longer deliberation times.
II.2.3 Evolution of asymmetry during the game
Beyond the final accumulated asymmetry, it is instructive to examine how the decision complexity imbalance develops move by move during the opening phase. Figure 6 presents the complexity asymmetry $A(n)=S_{B}(n)-S_{W}(n)$ as a function of ply $n$ (the number of moves made by each player) for all 960 Chess960 starting positions.
Figure 6: Evolution of the complexity asymmetry $A(n)=S_{B}(n)-S_{W}(n)$ as a function of ply $n$ for all 960 Chess960 starting positions. Gray lines show the trajectories of individual positions (each averaged over multiple games). The blue shaded region represents the ensemble mean (blue line) $\pm 1\sigma$. The green line with markers highlights position #518 (standard chess), with its corresponding shaded band indicating $\pm 1\sigma$. Parameters: $\Delta_{0}=10$ cp, depth 20, and $n=10$ moves per player.
At $n=0$, all positions begin with $A=0$ by definition—no moves have been made, so no decisions have yet been evaluated. As play progresses, the asymmetry trajectories diverge rapidly, forming a spreading fan of possibilities. By ply $n=10$, the ensemble spans a range of approximately [2.7,+1.8][-2.7,+1.8] bits, with mean $\langle A(10)\rangle=-0.37$ bits and standard deviation $\sigma_{A}=0.74$ bits. These values are consistent with the results shown in Fig. 4, although not identical due to the different engine depths used in the two analyses. The negative mean indicates that, on average across all Chess960 positions, White faces slightly more difficult decisions than Black during the opening phase.
The roughly linear decrease of the mean asymmetry with ply indicates that, on average, White faces increasingly harder decisions as the opening progresses. The widening spread shows that positions differ not only in the magnitude of their asymmetry but also in how quickly it develops: some configurations concentrate the complexity difference in the first few moves, while others accumulate it more gradually. The trajectories confirm that asymmetry is a dynamic property—some positions remain close to zero throughout, whereas others exhibit steadily growing imbalances. The progressively widening uncertainty band reflects both true variation across positions and game-to-game fluctuations for each configuration.
Standard chess (position #518) follows a distinctive trajectory that rises above the ensemble mean, reaching $\langle A(10)\rangle=+0.65\pm 1.57$ bits at the 92.1th percentile. This positive asymmetry indicates that Black accumulates greater decision complexity than White as the opening unfolds—a pattern that emerges consistently despite the large game-to-game variance. The trajectory’s upward drift becomes apparent by approximately ply 3–4, suggesting that the asymmetry in standard chess is not an artifact of the very first moves but rather a cumulative effect of the classical piece arrangement.
From a practical standpoint, these trajectories illuminate why certain Chess960 positions may feel more or less balanced during actual play. Positions with steep asymmetry growth place one player under cognitive pressure early in the game, potentially influencing time management and error rates. The observation that standard chess exhibits above-average asymmetry growth favoring Black’s burden aligns with the classical intuition that Black must work harder to equalize, though here we quantify this imbalance in information-theoretic terms rather than positional evaluation.
II.3 Key Positions in the Chess960 Complexity Landscape
II.3.1 Extreme positions and standard chess
The information cost analysis allows us to identify outlier configurations that lie at the extremes of the asymmetry and complexity spectra (Fig. 4). It should be noted that both the engine depth and the number of games used for averaging naturally influence these specific configurations; we present them here primarily to illustrate the broad diversity of situations that arise within the chess960 landscape.
The maximum asymmetry ‘favoring’ Black complexity is position #504 (RBQNBKNR): This configuration exhibits $\langle A\rangle=+1.84\pm 1.75$ bits, with Black experiencing $\langle S_{B}\rangle=5.47\pm 1.42$ bits compared to White’s $\langle S_{W}\rangle=3.63\pm 1.02$ bits. The back-rank arrangement places the rook on the usual a and h-files, bishops on b and e files flanking the queen on c, with the king on f. This seems to create a tactical environment where Black’s responses demand more evaluation depth than White’s developmental moves (the principal variation for white appears to be the quite natural 1.d4, 2.c4, 3.Nc3).
Conversely, the position #481 (BQRBNKNR) is the most complex for White: with $\langle A\rangle=-2.51\pm 1.65$ bits, this configuration places the heaviest burden on White ($\langle S_{W}\rangle=5.96\pm 1.35$ bits versus Black’s $\langle S_{B}\rangle=3.45\pm 0.94$ bits). The bishop on the a-file, queen on b, and rook on c create immediate tactical complications for White that constrain opening choices far more severely than Black’s (the main openings seem to revolve around a4, b4, c4, Ngf3 with no clear variation emerging).
The position #89 (NNRBBKRQ) achieves near-perfect symmetry (with the depth and number of games considered here) with $\langle A\rangle=-0.002\pm 1.65$ bits and $\langle S_{W}\rangle=4.47\pm 1.22$ bits, $\langle S_{B}\rangle=4.47\pm 1.11$ bits. This configuration, with knights on a and b files, rook on c, centrally paired bishops on d and e, and queen on h, creates a balanced tactical landscape where the decision complexity faced by both players is statistically indistinguishable. Note that even when the complexity is well balanced between Black and White, the Stockfish evaluation still favors White, with the principal variation beginning with one of the moves b4, Nb3, c4, ....
The ‘most complex game’, corresponding to the maximum total information cost, is obtained for position #226 (BNRQKBNR) with $\langle S_{\mathrm{tot}}\rangle=15.00\pm 1.02$ bits (99.9th percentile), with $\langle S_{W}\rangle=7.53\pm 0.56$ bits and $\langle S_{B}\rangle=7.47\pm 0.85$ bits. Remarkably, this configuration differs from the standard starting position by only a single transposition between the bishop and the rook on the a-file, yet it produces the highest overall decision complexity in the entire Chess960 ensemble. Notably, the position also exhibits near-perfect symmetry ($\langle A\rangle=-0.06$ bits), demonstrating that high complexity need not imply imbalanced cognitive demands. The relatively low standard deviations indicate that this complexity is consistently reproduced across independent game realizations.
The ‘simplest game’ is obtained for the position #316 (NBQRKRBN) with the smallest value $\langle S_{\mathrm{tot}}\rangle=4.89\pm 2.55$ bits (0th percentile), $\langle S_{W}\rangle=2.37\pm 1.83$ bits and $\langle S_{B}\rangle=2.52\pm 1.78$ bits. The queen on c-file between knight and rook, with the king centrally placed between rooks on d and f, probably limits early tactical complications. However, the large standard deviation indicates high variability—some games from this position can be considerably more complex than others.
@@ II.3.1 Extreme positions and standard chess
Standard chess (position #518, RNBQKBNR) exhibits $\langle S_{W}\rangle=4.29\pm 1.22$ bits and $\langle S_{B}\rangle=4.97\pm 1.17$ bits, yielding asymmetry $\langle A\rangle=+0.69\pm 1.69$ bits and total complexity $\langle S_{\mathrm{tot}}\rangle=9.26\pm 1.69$ bits. The positive asymmetry places standard chess at the 91th percentile, indicating that Black faces moderately harder decisions than in most Chess960 configurations. However, the uncertainty is large: zero asymmetry lies within one standard deviation, so this imbalance is not statistically significant at the single-game level. The total complexity percentile of 47th indicates that standard chess exhibits moderate overall complexity—neither exceptionally simple nor complex relative to the broader Chess960 ensemble.
@@
We illustrate the large variety of both complexity and evaluation in the Fig. 7 that presents both the complexity asymmetry $A$ and the initial Stockfish evaluation $E$ as functions of the Chess960 position number. This visualization highlights the substantial diversity of both complexity and initial advantage across the full ensemble of 960 starting configurations.
Figure 7: Complexity asymmetry $A(n)=S_{B}(n)-S_{W}(n)$ for $n=10$ plies (top) and initial Stockfish evaluation $E$ (bottom) as functions of Chess960 position number. The shaded region in panel (a) represents $\pm 1\sigma$ uncertainty from multi-game averaging. Gold stars mark position #518 (standard chess). Red dashed lines indicate ensemble means ($\langle A\rangle$ and $\langle E\rangle$). The position numbering follows the standard Chess960 enumeration scheme based on back-rank piece placement. Both quantities show substantial variation across positions with no obvious structure. Calculations were done at depth 14, averaged over 5 games and $\Delta_{0}=10$cp.
The asymmetry panel (Fig. 7a) displays considerable variation across positions, with values spanning approximately [2.51,+1.84][-2.51,+1.84] bits around a slightly negative mean. The shaded uncertainty band illustrates the game-to-game variability for each position; notably, this variability is comparable in magnitude to the inter-position differences, underscoring the stochastic nature of individual game trajectories. Position #518 (standard chess) lies slightly above the mean asymmetry, consistent with its 91st percentile ranking.
The evaluation panel (Fig. 7b) shows that initial Stockfish assessments cluster tightly around the mean, with most positions evaluated within $\pm 0.1$ pawns. This narrow distribution reflects the fundamental balance built into Chess960’s design constraints: all positions satisfy the same castling prerequisites and bishop color requirements, limiting gross structural imbalances. The few outlier evaluations correspond likely to positions where piece placement creates immediate tactical asymmetries.
Importantly, the two panels reveal no obvious correlated structure: positions with extreme asymmetry $A$ do not systematically coincide with extreme evaluations $E$, and vice versa. The position numbering scheme, based on the Scharnagl encoding of back-rank arrangements, could introduce some local correlations—neighboring positions often share similar piece placements—but no clear correlations or periodic structure emerges in either quantity. This likely reflects the strong sensitivity of chess to microscopic details: a change as small as moving one piece by a single square can redirect the entire course of the game, leading to dramatically different evaluations and strategic outcomes.
From a game-design perspective, these findings illuminate the properties of the classical starting position: standard chess maintains moderate complexity (47th percentile) while exhibiting above-average asymmetry favoring Black’s decision burden (91st percentile). Position #226 (BNRQKBNR) offers an interesting alternative: it achieves the highest complexity in the ensemble while maintaining near-perfect symmetry, potentially providing both maximal strategic depth and balanced competitive conditions.
II.3.2 Looking for the most balanced position
@@
The two fairness metrics introduced above—initial evaluation $E$ and decision asymmetry AA—capture distinct aspects of competitive balance: the former measures the expected outcome advantage, while the latter quantifies the imbalance in cognitive burden between players. A natural question is whether these metrics are correlated: do positions favoring White in evaluation also impose harder decisions on Black?
@@
Figure 8 plots the complexity asymmetry $A=S_{B}-S_{W}$ against the initial Stockfish evaluation $E$ for all 960 starting positions. The correlation is weak ($r=0.15$), indicating that evaluation advantage and decision complexity are largely independent dimensions of positional character. Positions favoring White in evaluation do not systematically impose harder decisions on Black, nor do evaluation-balanced positions guarantee symmetric cognitive demands.
@@
Figure 8: Complexity asymmetry $A=S_{B}-S_{W}$ versus initial Stockfish evaluation $E$ (depth 30) for all 960 Chess960 starting positions. Error bars indicate one standard deviation across 5 games per position. The weak correlation ($r=0.15$) shows that evaluation balance and decision symmetry are largely independent. The gold star marks position #518 (standard chess); the green diamond indicates position #198 (QNBRKBNR), the most balanced configuration minimizing distance to the origin in the normalized $(E,A)$ plane.
This independence motivates a search for the configuration that minimizes both imbalances simultaneously. We define the most balanced position as the one minimizing the normalized distance to the origin
$d=\sqrt{\left(\frac{E}{\sigma_{E}}\right)^{2}+\left(\frac{A}{\sigma_{A}}\right)^{2}}$ (7)
where $\sigma_{E}$ and $\sigma_{A}$ are the standard deviations of evaluation and asymmetry across the ensemble, respectively. This criterion identifies position #198 (QNBRKBNR) as the most balanced configuration, with $E=+0.03$ pawns and $A=-0.03\pm 1.68$ bits—both effectively zero within measurement uncertainty. Surprisingly, this arrangement can be obtained from the standard chess position by a single transposition between the queen and the rook on the a-file. Because the initial evaluation is very small, many openings perform similarly well, including 1.b3, 1.d4, and 1.c4.
Standard chess (position #518), by contrast, lies far from this optimum: while its evaluation is typical ($E=+0.30$ pawns), its decision asymmetry places it at the 91st percentile ($A=+0.69\pm 1.69$ bits), indicating that Black faces a greater cognitive burden than in most Chess960 configurations. The existence of configurations like position #198 that achieve near-perfect balance in both metrics suggests that tournament organizers seeking maximally fair starting positions have principled alternatives to both standard chess and random selection.
Clearly, further chess-specific analysis is required to understand why this position exhibits such balanced behavior (and why the ‘extreme’ positions discussed above have their speific properties). Relevant features may include the placement of the rooks (for example, whether both are on the queenside), the king’s position and the resulting difficulty or subtlety of castling decisions, the presence of early tactical opportunities that must be avoided, the initial geometry of the bishops—whether they immediately control long diagonals and if they operate on the same flank—and the availability of viable pawn breaks in the opening. A more detailed examination of these structural elements is needed to translate the statistical findings into concrete chess insights.
III Discussion
Our study identifies two main insights linking chess to broader ideas in complex systems: the 960 starting positions form a rugged landscape of decision complexity, and small structural changes can significantly alter the balance and difficulty of the game.
First, standard chess (position #518) occupies a statistically unremarkable location in the Chess960 complexity landscape: it shows a typical initial advantage ($E=+0.30$ pawns) and a moderate total complexity (47th percentile), while its asymmetry lies in the upper tail of the distribution (91st percentile), indicating that Black faces somewhat harder decisions in the opening. This challenges the intuitive notion that centuries of play have optimized the classical starting position for maximal strategic depth. Several hypotheses may account for this observation. The classical arrangement exhibits high visual symmetry about the board’s vertical axis, making it both aesthetically pleasing and easy to remember—a practical advantage for cultural transmission. Additionally, moderate complexity may be optimal for learning, as overly simple positions lack strategic richness while highly complex ones prove too chaotic for systematic study; the classical position may thus represent an implicit compromise between depth and accessibility. Finally, it is possible that no optimization occurred at all—the classical configuration could simply reflect historical accident, frozen by tradition once chess rules stabilized in the 15th century.
Second, we find substantial heterogeneity across the Chess960 ensemble in both complexity and fairness. The asymmetry parameter $A=S_{b}-S_{w}$ spans more than 4 bits, with some positions strongly favoring White and others Black in terms of decision-making burden. This heterogeneity has practical implications: in Freestyle chess tournaments, starting positions are typically drawn at random under the assumption that all configurations are roughly equivalent. Our results challenge this assumption—random selection does not guarantee competitive fairness, as some positions impose significantly unequal cognitive demands on the two players. This finding provides a quantitative basis for more principled position selection protocols in competitive Chess960, whether by excluding outliers in asymmetry, curating balanced subsets, or weighting draws toward fairer configurations.
IV Conclusion
In this work, we introduced an information–cost framework for quantifying decision complexity in Chess960 starting positions. Our analysis shows that the 960 configurations differ markedly in both strategic depth and competitive fairness, challenging the common assumption that all starting positions are equivalent. Standard chess, despite centuries of cultural evolution, does not occupy an exceptional location in this landscape: it exhibits a typical initial advantage and moderate total complexity, while displaying above-average asymmetry in decision difficulty. This may reflect implicit historical selection for positions that balance richness with accessibility rather than maximal complexity or perfect fairness.
More broadly, our results demonstrate how concepts from information theory and statistical physics can quantify strategic complexity in deterministic decision-making systems. The framework presented here is not specific to chess; it applies to any setting in which optimal choices can be evaluated and the difficulty of distinguishing them can be measured.
These findings also open perspectives for studying the cultural evolution of board games. The fact that the classical starting position is not structurally special raises questions about the ‘phylogenetics’ of chess variants: how the traditional configuration emerged from the broader space of possibilities, and what selection pressures—cognitive, aesthetic, pedagogical, or historical—shaped its adoption. Our approach provides quantitative tools for comparing historical games (Shatranj, Xiangqi, Shogi) and for reconstructing the underlying fitness landscape that may have influenced their evolution. It also allows one to examine whether successful games cluster in particular regions of complexity space, and whether convergent evolution toward similar complexity profiles occurred independently across cultures.
Finally, the framework can be extended by incorporating higher-order information-theoretic measures—such as mutual information between successive moves—to probe the dynamical structure of openings. Applying similar methods to other randomized-start variants or to entirely different strategic games (Go, Shogi, and modern board games) would test the generality of these ideas and enable systematic cross-game comparisons of strategic complexity.

### v3

@@ II.4 Characterizing the opening complexity
which quantifies the differential complexity: positive values indicate Black faces more difficult decisions, while negative values indicate White’s burden is greater. Analysis of all 960 starting positions (Fig. 4(d)) reveals a mean asymmetry of $\langle A\rangle=-0.26$ bits (and standard deviation $\pm 1.36$ bits), suggesting a weak but systematic bias wherein White generally faces slightly more complex decision trees. This negative mean may reflect the fundamental asymmetry inherent in chess: White moves first and must immediately confront the full complexity of opening choices, while Black’s initial responses are partially constrained by White’s play. The substantial variation from 4.4-4.4 to +4.2+4.2 bits) indicates high game-to-game variability, reflecting the sensitivity of decision complexity to openings for specific initial conditions.
@@
Standard chess (position #518, RNBQKBNR) exhibits $\langle S_{W}\rangle=5.42$ bits and $\langle S_{B}\rangle=5.79$ bits, yielding asymmetry $\langle A\rangle=+0.36$ bits and total complexity $\langle S_{\mathrm{tot}}\rangle=11.2$ bits. The positive asymmetry places standard chess at the 69th percentile, indicating that Black faces moderately harder decisions than in most Chess960 configurations. The total complexity percentile of 72.5th indicates that standard chess exhibits above average overall complexity, but can be considered as neither exceptionally simple nor complex relative to the broader Chess960 ensemble.
@@
II.5 Illustration 1: Searching for jointly balanced positions
@@
The two fairness metrics introduced above—initial evaluation $E$ and decision asymmetry AA—capture distinct aspects of competitive balance: $E$ measures the expected outcome advantage, while $A$ quantifies the imbalance in cognitive burden between players. A natural question is whether these two dimensions are correlated. Do positions favoring White in evaluation also impose greater decision complexity on Black?
@@
Figure 5 plots the asymmetry $A=S_{B}-S_{W}$ against the initial Stockfish evaluation $E$ (depth 15) for all 960 starting positions. The correlation is weak ($r\approx 0.15$), indicating that evaluation advantage and decision complexity are largely independent characteristics of a position. Positions that are evaluation-balanced do not necessarily yield balanced cognitive demands, and vice versa.
@@
Figure 5: Complexity asymmetry $A=S_{B}-S_{W}$ versus initial Stockfish evaluation $E$ (Stockfish 17.1, depth 15, single thread) for all 960 Chess960 starting positions. The weak correlation ($r\approx 0.15$) indicates near-independence between evaluation balance and decision symmetry. The gold star marks position #518 (standard chess). The green diamond indicates the configuration that minimizes the normalized distance to the origin in the $(E,A)$ plane at this depth.
This near-independence motivates the search for configurations that are simultaneously close to zero in both metrics. We therefore define a joint balance score as the normalized distance
$d=\sqrt{\left(\frac{E}{\sigma_{E}}\right)^{2}+\left(\frac{A}{\sigma_{A}}\right)^{2}},$ (5)
where $\sigma_{E}$ and $\sigma_{A}$ are the ensemble standard deviations of $E$ and $A$, respectively. This provides a scale-invariant measure of proximity to the ideal balanced point $(E,A)=(0,0)$.
At depth 15, this criterion identifies position #823 (RKBNQRNB) as the closest configuration to the origin, with $E\approx 0$ pawns and $A\approx+0.26$ bits. Importantly, however, the identity of the minimizing position is not universal: as discussed above, outliers and fine rankings can depend on engine depth and evaluation settings. The purpose of this construction is therefore not to single out a unique “optimal” starting position, but rather to illustrate how joint fairness criteria can be defined and systematically explored.
Standard chess (position #518) lies at a moderate distance from the origin, and although standard chess is not extreme in evaluation, it exhibits a non-negligible structural decision asymmetry.
A deeper chess-specific interpretation of jointly balanced configurations requires further analysis. Relevant structural factors may include rook placement (e.g., both rooks on the same flank), king position and castling structure, early tactical constraints, bishop geometry and diagonal accessibility, and the availability of immediate pawn breaks. Understanding how such geometric features influence both $E$ and $A$ would provide a bridge between statistical characterization and concrete opening theory.
II.6 Illustration 2: the most complex positions
With this consolidated protocol (and depth $15$), the most complex Chess960 initial position is position $524$ (RBNQKNBR), with a total complexity $S_{\mathrm{tot}}=17.17$ bits. For comparison, the standard chess starting position (position $518$, RNBQKBNR) has $S_{\mathrm{tot}}=11.20$ bits (and initial evaluation $E=39$ cp).
Interestingly, this most complex position remains very close to the standard chess configuration, differing only by the permutation of a few pieces (B and N on both king and queen sides). Its engine evaluation is also relatively small, $E=24$ cp, which is even closer to equality than the standard chess position. This observation confirms that high combinatorial complexity is not associated with a large initial imbalance in the engine evaluation.
The other configurations with the largest complexity have also values far above that of the standard starting position. The five most complex positions identified in our analysis are: position $524$ (RBNQKNBR) with $S_{\mathrm{tot}}=17.166$ ($S_{W}=8.886$, $S_{B}=8.280$); position $444$ (RBNNQKBR) with $S_{\mathrm{tot}}=16.685$ ($S_{W}=8.190$, $S_{B}=8.495$); position $932$ (RBBKRNQN) with $S_{\mathrm{tot}}=15.746$ ($S_{W}=7.383$, $S_{B}=8.362$); position $535$ (RNBKQNRB) with $S_{\mathrm{tot}}=15.540$ ($S_{W}=7.756$, $S_{B}=7.784$); position $372$ (NBBRKRNQ) with $S_{\mathrm{tot}}=15.277$ ($S_{W}=7.896$, $S_{B}=7.381$).
It would be interesting to understand what these positions have in common and what governs their large complexity values. Finally, it is important to stress that these values are obtained at search depth $15$. While this depth already provides a reasonable proxy for human-level analysis, deeper searches may modify the ranking of the most complex positions, as discussed above. The configurations reported here should therefore be interpreted as belonging to a family of extremely high-complexity starting positions rather than as defining an absolutely fixed ordering across all possible search depths.
III Discussion and Conclusion
We introduced an information–cost framework to quantify decision complexity across all 960 Chess960 starting positions. Two main results emerge. First, the Chess960 ensemble forms a highly heterogeneous landscape: total opening complexity and decision asymmetry vary widely, showing that small rearrangements of back-rank pieces can substantially alter both strategic depth and competitive fairness. In particular, the asymmetry $A=S_{B}-S_{W}$ spans several bits, with some positions disproportionately burdening White and others Black. A small random sample of starting positions does not, in general, guarantee competitive balance. Our framework provides a quantitative basis for designing more principled selection protocols in Freestyle tournaments.
Second, the classical starting position (position #518) occupies a statistically unremarkable location in this landscape. It exhibits a typical evaluation advantage for White, moderate total complexity, and a slightly positive asymmetry indicating a greater cognitive burden for Black. Despite centuries of play, the traditional arrangement does not maximize complexity nor minimize imbalance. This suggests that its historical persistence may reflect aesthetic symmetry, pedagogical accessibility, or cultural contingency rather than structural optimality.
More broadly, this work demonstrates how tools from information theory and statistical physics can characterize strategic complexity in deterministic decision systems. The framework is not specific to chess: it applies whenever optimal actions can be evaluated and the informational effort required to distinguish them can be quantified. Extending this approach to other chess variants, historical games, or different strategic systems (Go, Shogi, and modern board games) would allow systematic comparisons of complexity landscapes and provide new insight into the evolution and design of competitive games.